OpenStudy (anonymous):

square root of (-72) then all ^3

OpenStudy (anonymous):

\[\sqrt{-72}^3\]

OpenStudy (anonymous):

sqrt(-72)^3 = - 610.940259 i

OpenStudy (anonymous):

Think of it as -72^(3/2), because square root is just x^(1/2), and 3 * (1/2) = 3/2. So (-72)^(3/2)